OpenStudy (anonymous):

For y = x2 − 4x + 3, Determine if the parabola opens up or down. State if the vertex will be a maximum or minimum. Find the vertex. Find the x-intercepts. Describe the graph of the equation. Show all work and use complete sentences

hero (hero):

On the graph of y = x^2 - 4x + 3, the parabola opens upward. The vertex will therefore be a minimum. To find the vertex, use x = -b/2a, then substitute x back into the original equation to find y. To find the x-intercept, set y = 0, then solve for x, The graph of y = x^2 - 4x + 3 is the graph of x^2 shifted two units to the right, and 1 unit downward.